OpenStudy (anonymous):

please help me slove these problems 1.) 6xy=42 2.) x-1/2y=-1 3.)2y+x=8 4.) 7=1/3y+1 5.) F=9/5 C+32

OpenStudy (alienium):

geez... don't ask for your homeworks before trying to solve them by your own

OpenStudy (anonymous):

i have

OpenStudy (alienium):

1. has a range of possible solutions 2. x-1/2y = -1 (x-1/2y) + 1 = 0 2xy-2y+2y = 0 2xy = 0 x or y = 0 2. 2y+x = 8 y+x= 4 (range of solutions) 3. 7 = 1/3y+1 (1/3y+1) - 7 = 0 3y+1-21y+7=0 -18y+8=0 y = 8/-18 4. i don't understand

OpenStudy (anonymous):

wait you set it =to 0?

OpenStudy (alienium):

yes... if you solve the base case it's simple just pass the value in the other side and add it to the value of x

OpenStudy (anonymous):

all of these are incorrect. i have a sheet where its like cross word none of these answers are there..thanks any way.